# Bayesian Estimation and Tuning-Free Rank Detection for Probability Mass Function Tensors

### v1

@@ Abstract
Obtaining a reliable estimate of the joint probability mass function (PMF) of a set of random variables from observed data is a significant objective in statistical signal processing and machine learning. Modelling the joint PMF as a tensor that admits a low-rank canonical polyadic decomposition (CPD) has enabled the development of efficient PMF estimation algorithms. However, these algorithms require the rank (model order) of the tensor to be specified beforehand. In real-world applications, the true rank is unknown. Therefore, an appropriate rank is usually selected from a candidate set either by observing validation errors or by computing various likelihood-based information criteria, a procedure which is computationally expensive for large datasets. This paper presents a novel Bayesian framework for estimating the joint PMF and automatically inferring its rank from observed data. We specify a Bayesian PMF estimation model and employ appropriate prior distributions for the model parameters, allowing for tuning-free rank inference via a single training run. We then derive a deterministic solution based on variational inference (VI) to approximate the posterior distributions of various model parameters. Additionally, we develop a scalable version of the VI-based approach by leveraging stochastic variational inference (SVI) to arrive at an efficient algorithm whose complexity scales sublinearly with the size of the dataset. Numerical experiments involving both synthetic data and real movie recommendation data illustrate the advantages of our VI and SVI-based methods in terms of estimation accuracy, automatic rank detection, and computational efficiency.
Index Terms:
Statistical learning, Bayesian inference, tensor decomposition, joint PMF estimation, rank detection, recommender systems
@@ I Introduction
In statistical signal processing and machine learning, one of the most essential, yet challenging, problems is obtaining a reliable estimate of the joint distribution of a set of random variables from observed data. In the case of multivariate discrete random variables, the joint distribution is represented by a joint probability mass function (PMF) which describes the probability of each of the possible realizations of the random variables. Knowledge of the joint PMF enables many statistical learning tasks to be solved in an explainable and optimal manner. For instance, two common tasks which are ubiquitous in statistical learning are predicting missing data (e.g., item ratings) given a subset of observations and predicting class/cluster labels given the associated features. Applications of these learning tasks arise in recommender systems (e.g., for movies, music, etc.) and data classification/clustering, respectively. Using the joint PMF, the posterior distribution of the missing data given the observed data can be computed, from which the conditional expectation of missing data can be found. Similarly, the posterior distribution of the class labels given the features can be computed from the joint PMF. The predicted class label is then the maximizer of the posterior distribution. It is known that these estimates are optimal in that they minimize the mean squared error and the probability of misclassification, respectively (see, e.g., [1]).
@@
A classical non-parametric technique for estimating a joint PMF from observed data is histogram estimation, i.e., counting the number of instances each possible realization of the random variables is observed. However, this approach faces the curse of dimensionality: the number of observations required for a reliable estimate grows exponentially with the number of variables. This drawback limits the applicability of histogram estimation for many practical cases of interest. Therefore, recent work on PMF estimation has focused on developing alternative approaches which mitigate the curse of dimensionality. These approaches exploit the fact that a joint PMF of a set of $N$ random variables can be represented as an $N$-dimensional tensor which admits a low-rank canonical polyadic decomposition (CPD) (see, e.g., [2], [3]). Importantly, the work in [4] showed the connection between various latent variable models and tensor decompositions. In particular, the author drew a parallel between the naïve Bayes model and the CPD. This connection was proved more rigorously in [5], where it was shown that any joint PMF can be represented by a naïve Bayes model with one latent variable taking a finite number of states. Furthermore, the number of latent states is the rank of the joint PMF tensor, i.e., the minimum number of rank-one components in the CPD. The authors then developed a PMF estimation algorithm based on coupled tensor factorization of third-order marginal PMF tensors. Further works have proposed various PMF estimation approaches, ranging from coupled tensor factorization based on lower-order marginals [6, 7, 8, 9, 10] to maximum-likelihood based approaches in which no lower-order marginals need to be computed [11, 12].
@@
In all the aforementioned approaches, the rank (or the model order) of the joint PMF tensor needs to be somehow specified before estimation. However, while there exist some results regarding bounds on the tensor rank (see, e.g., [2], [3]), in real-world scenarios, the exact tensor rank is unknown. In practice, the most common approach is to treat the rank as a hyperparameter to be tuned via cross-validation (CV) (this is done, e.g., in [5]). Alternatively, there are well-established model order selection criteria based on likelihood functions. These include the Akaike information criterion (AIC) [13] and the Bayesian information criterion (BIC) [14]. In recent years, the decomposed normalized maximum likelihood criterion (DNML) [15], which is based on the minimum description length principle, has also been proposed for model order selection in hierarchical latent variable models. While these techniques are well-established and have theoretical guarantees, they require that the rank be selected from a prespecified candidate set. For PMF estimation problems involving large datasets, this model order selection procedure could be costly in terms of computation time. One way to curb the high computation cost could be to decrease the size of the candidate set, e.g., by decreasing the range of values or by having larger intervals between the elements. However, there is always the possibility of model order misspecification in case the true model order is not included in the candidate set.
Automatic methods for determining the rank of a tensor have been extensively studied in the context of probabilistic tensor factorization [16, 17, 18, 19]. In these works, the goal is to obtain a low-rank decomposition of tensorized (i.e., multidimensional) data and to simultaneously infer the rank of the tensor. This is typically achieved by assigning sparsity-promoting prior distributions which constrain the number of low-rank components to a minimum, allowing the rank to be determined as part of the inference procedure. For tractability, posterior inference is carried out using variational inference (VI) [20] which is a well-established method for approximating posterior distributions.
In this paper, we propose a novel Bayesian framework to estimate the joint PMF tensor of a set of discrete random variables and automatically infer the rank of the tensor from observed data. In contrast to existing PMF estimation techniques, our approach requires no parameter tuning, thus paving the way for efficient PMF estimation for large-scale problems due to the reduction in computational complexity. Our work has the following specific contributions:
  • We propose a deterministic Bayesian algorithm for joint PMF tensor estimation and automatic rank inference. We exploit the naïve Bayes representation of the joint PMF to specify a probabilistic PMF estimation model. In particular, we assign Dirichlet priors to the CPD components (the model parameters), namely, the factor matrices and the loading vector. Due to the underlying connection between the low-rank model and the naïve Bayes model, this choice of a prior distribution is important since the CPD components are constrained to the probability simplex (i.e., nonnegativity and sum-to-one). In addition, we use a sparsity-promoting Dirichlet prior for the loading vector, which controls the number of rank-one CPD components, allowing for automatic rank detection. Our choice of prior leads to a Bayesian model which is markedly different from the existing models used in probabilistic tensor factorization, e.g., [16, 17, 18, 19]. Performing exact Bayesian inference to determine the posterior distribution of the model parameters given the data is computationally infeasible with our resulting model. Therefore, we invoke VI to derive a deterministic solution that approximates the posterior distributions of the model parameters. An initial version of the VI-based approach was presented at CAMSAP 2023 [21]. Here, we provide more details on the Bayesian model specification and the derivation of important expressions, an in-depth discussion on the automatic rank inference properties of the algorithm, as well as an appendix elaborating how the optimization objective function is derived. Moreover, in [21], the algorithm was only tested on synthetic data. Therefore, in this work, we present results based on real data to further validate our method.
  • We propose a scalable Bayesian PMF estimation algorithm by extending the deterministic VI-based approach. The computational complexity of the VI-based algorithm grows linearly with the size of the dataset. While this is tolerable for small and medium-sized problems, it is desirable to have an algorithm which can scale to large problem sizes involving many random variables and/or massive datasets. To this end, we develop an efficient estimation algorithm whose computational complexity scales sub-linearly with the size of the dataset. The algorithm is based on stochastic variational inference (SVI), a scalable VI-based method that was originally applied to topic modeling [22]. Recently, SVI has been applied to biomedical data analysis [23] and probabilistic tensor factorization [24, 25]. To the best of our knowledge, this work is the first to present an SVI-based algorithm for joint PMF tensor estimation and rank detection. The algorithm computes noisy gradients of the objective function from randomly sampled minibatches of data, in contrast to the VI-based approach, where the entire dataset needs to be analyzed in each iteration before the posterior distributions can be updated. This SVI approach presents remarkable gains in computational efficiency, making it suitable for big data applications.
  • We validate our Bayesian PMF estimation approaches based on extensive numerical simulations. Experiments on synthetic data demonstrate the advantages of our methods in terms of automatic rank detection, robustness to overfitting, and computational efficiency. Additionally, results from a real-world movie recommendation experiment reveal that our methods outperform state-of-the-art approaches in terms of accuracy and computational efficiency.
I-A Notation
@@
This paper uses bold lowercase letters (e.g., $\bm{a}$) to denote vectors, bold uppercase letters (e.g., $\bm{A}$) to represent matrices, and bold calligraphic letters (e.g., $\bm{\mathcal{A}}$) to represent tensors. Uppercase letters (e.g., $A$) denote scalar random variables while lowercase letters (e.g. $a$) denote realizations of the corresponding random variables. The outer product is represented by $\circ$ while T denotes the transpose operator.
@@ II Problem Statement
Consider a discrete random vector $\bm{x}=[X_{1},\dotsc,X_{N}]^{\mathsf{T}}\in\mathbb{R}^{N}$ where each random variable $X_{n}$ can take discrete values in $\{1,\dotsc,I_{n}\}$, $n=1,\dotsc,N$. The joint PMF $p(\bm{x})$ can be described by an $N$-dimensional tensor $\bm{\mathcal{X}}\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$ whereby each element of the tensor is the joint probability of a particular realization of the random variables, i.e., $\mathsf{Pr}(X_{1}=i_{1},\dotsc,X_{N}=i_{N})=\bm{\mathcal{X}}(i_{1},\dotsc,i_{N})$. In general, a tensor $\bm{\mathcal{X}}$ can be expressed as a sum of rank-one terms via the CPD
@@
where $\bm{A}_{n}\in\mathbb{R}^{I_{n}\times R}$ denotes a factor matrix and $\bm{A}_{n}(:,r)$ denotes the $r$-th column of $\bm{A}_{n}$. Without loss of generality, the columns of the factor matrices can be restricted to have unit norm such that $\|\bm{A}_{n}(:,r)\|_{p}=1$ for $p\geq 1$, $\forall~{}n,r$. Then, $\bm{\lambda}=[\lambda_{1},\dotsc,\lambda_{r}]^{\mathsf{T}}$ is the loading vector which ‘absorbs’ the norms of the columns. The minimum number $R$ of rank-1 terms required for the decomposition to hold is referred to as the rank of the tensor. In terms of the CPD, a particular element of the tensor is given by
@@
It has been shown in [4] and [5] that a rank-$R$ CPD can be interpreted as a particular kind of latent variable model known as the naïve Bayes model. This model assumes that the random variables $\{X_{n}\}_{n=1}^{N}$ are conditionally independent given a hidden (latent) variable $H$ which can take a finite number of $R$ states. Under this interpretation, each element $\lambda_{r}$ of the loading vector is the probability $\mathsf{Pr}(H=r)$ of the hidden variable, while each factor matrix column $\bm{A}_{n}(:,r)$ is the conditional PMF $p(X_{n}\,|\,H=r)$. Thus, (1) is subject to a set of probability simplex constraints, i.e., $\bm{\lambda}>\bm{0},\bm{A}_{n}\geq\bm{0}$ (nonnegativity) and $\bm{1}^{\mathsf{T}}\bm{\lambda}=1$, $\bm{1}^{\mathsf{T}}\bm{A}_{n}=\bm{1}^{\mathsf{T}}$ (sum-to-one).
@@
(independently for each $n$ and independently of the values in $\bm{x}$), where the outage probability $p$ denotes the probability that $X_{n}$ is unobserved in $\bm{y}$. We have at hand a dataset $\bm{Y}=[\bm{y}_{1},\dotsc,\bm{y}_{T}]$ containing $T$ i.i.d. realizations of $\bm{y}$, where $\bm{y}_{t}=[y_{1,t},\dotsc,y_{N,t}]^{\mathsf{T}}$. The goal is to estimate the joint PMF tensor $\bm{\mathcal{X}}$ from the observations $\bm{Y}$.

### v2

# Joint Bayesian Parameter and Model Order Estimation for Low-Rank Probability Mass Tensors

@@ Abstract
Obtaining a reliable estimate of the joint probability mass function (PMF) of a set of random variables from observed data is a significant objective in statistical signal processing and machine learning. Modelling the joint PMF as a tensor that admits a low-rank canonical polyadic decomposition (CPD) has enabled the development of efficient PMF estimation algorithms. However, these algorithms require the rank (model order) of the tensor to be specified beforehand. In real-world applications, the true rank is unknown. Therefore, an appropriate rank is usually selected from a candidate set either by observing validation errors or by computing various likelihood-based information criteria, a procedure that could be costly in terms of computational time or hardware resources, or could result in mismatched models which affect the model accuracy. This paper presents a novel Bayesian framework for estimating the low-rank components of a joint PMF tensor and simultaneously inferring its rank from the observed data. We specify a Bayesian PMF estimation model and employ appropriate prior distributions for the model parameters, allowing the rank to be inferred without cross-validation. We then derive a deterministic solution based on variational inference (VI) to approximate the posterior distributions of various model parameters. Numerical experiments involving both synthetic data and real classification and item recommendation data illustrate the advantages of our VI-based method in terms of estimation accuracy, automatic rank detection, and computational efficiency.
@@ I Introduction
In statistical signal processing and machine learning, one of the most essential, yet challenging, problems is obtaining a reliable estimate of the joint distribution of a set of random variables from observed data. In the case of multivariate discrete random variables, the joint distribution is represented by a joint probability mass function (PMF) which describes the probability of each of the possible realizations of the random variables. Knowledge of the joint PMF enables many statistical learning tasks to be solved in an explainable and optimal manner. For instance, two common tasks which are ubiquitous in statistical learning are predicting missing data (e.g., item ratings) given a subset of observations , and predicting class or cluster labels given the associated features [1]. Applications of these learning tasks arise in recommender systems (e.g., for movies, music, etc.) and data classification or clustering, respectively [2, 3]. Using the joint PMF, the posterior distribution of the missing data given the observed data can be computed, from which the conditional expectation of missing data can be found. Similarly, the posterior distribution of the class labels given the features can be computed from the joint PMF. The predicted class label is then the maximizer of the posterior distribution. It is known that these estimates are optimal in that they minimize the mean squared error and the probability of misclassification, respectively (see, e.g., [3]).
@@
A classical nonparametric technique for estimating a joint PMF from observed data is histogram estimation, i.e., counting the relative number of instances each possible realization of the random variables is observed. However, this approach faces the curse of dimensionality: the number of observations required for a reliable estimate grows exponentially with the number of variables. This drawback limits the applicability of histogram estimation for many practical cases of interest. Therefore, recent work on PMF estimation has focused on developing alternative approaches which mitigate the curse of dimensionality. These approaches exploit the fact that a joint PMF of a set of $N$ random variables can often be represented as an $N$-dimensional tensor which admits a low-rank canonical polyadic decomposition (CPD) (see, e.g., [4], [5]). Importantly, the work in [6] showed the connection between various latent variable models and tensor decompositions. In particular, the author drew a parallel between the naïve Bayes model and the CPD. This connection was proved more rigorously in [7], where it was shown that any joint PMF can be represented by a naïve Bayes model with one latent variable taking a finite number of states. Furthermore, the number of latent states is the rank of the joint PMF tensor, i.e., the minimum number of rank-one components in the CPD. In addition, the authors proved that the joint PMF of all the variables can be recovered from lower-order (e.g., third-order) marginal PMF tensors. Further works have proposed various PMF estimation approaches, ranging from coupled tensor factorization based on lower-order marginals [8, 9, 10, 11] to maximum-likelihood based approaches in which no lower-order marginals need to be computed [12, 13].
@@
In all the aforementioned approaches, the rank (or the model order) of the joint PMF tensor needs to be somehow specified before estimation. However, while there exist some results regarding bounds on the tensor rank (see, e.g., [4], [5]), in real-world scenarios, the exact tensor rank is unknown. In practice, the most common approach is to treat the rank as a hyperparameter to be tuned via cross-validation (CV) (this is done, e.g., in [7]). Alternatively, there are classical model order selection criteria based on likelihood functions. These include the Akaike information criterion (AIC) [14] and the Bayesian information criterion (BIC) [15]. In recent years, the decomposed normalized maximum likelihood criterion (DNML) [16], which is based on the minimum description length principle, has also been proposed for model order selection in hierarchical latent variable models. While these techniques are well-established and have theoretical guarantees, they require that the rank be selected from a prespecified candidate set. If the training cost of the estimation algorithm is high (e.g., for datasets with many samples or random variables), such a model order selection procedure is costly in terms of computation time or hardware resources due to the requirement to carry out multiple training runs. One way to reduce the computation cost could be to decrease the size of the candidate set, e.g., by limiting the range of tested orders or by testing orders on a sparse grid. However, there is always the possibility of model order misspecification in case the true model order is not included in the candidate set.
The objective of this paper is to develop an efficient method to estimate the parameters of a low-rank joint PMF tensor as well as its model order, or rank. Our contributions are summarized as follows:
  • We propose a novel and efficient Bayesian framework to estimate both the low-rank joint PMF tensor of a set of discrete random variables and its rank in a unified manner. Our method enables the rank of the PMF tensor to be estimated as part of the Bayesian inference procedure without resorting to cross-validation or other model order selection techniques as in existing PMF tensor estimation approaches (Section III).
  • The connection between the naïve Bayes model and the CPD of a joint PMF tensor is exploited to specify a Bayesian model, tailored to low-rank PMF estimation, that ensures that the estimated parameters satisfy the probability simplex (nonnegativity and sum-to-one) constraints. A sparsity-promoting Dirichlet prior is imposed on the loading vector (the weights of the latent components) to enable identification and pruning of irrelevant components without heuristic thresholding or manual model selection (Subsection III-A).
  • We develop a PMF estimation algorithm based on variational inference (VI) to approximate posterior distributions over the CPD model parameters. The VI approach provides closed-form updates for all posterior distributions, enabling efficient implementation and deterministic convergence without sampling. Moreover, we derive a pruning threshold for irrelevant components that depends on the prior hyperparameter and the dataset size, avoiding ad hoc threshold specification (Subsections III-B, III-D).
  • We validate the accuracy and rank estimation performance of our proposed method through carefully designed synthetic data simulations, which include varying dataset sizes, PMF tensor ranks, outage probabilities (fraction of missing data), and robustness to the choice of the hyperparameter for the loading vector prior (Section IV). Moreover, we present results using real datasets to demonstrate applicability to practical classification and item recommendation tasks (Section V).
An initial version of the VI-based approach was presented at CAMSAP 2023 [17]. Here, we provide more details on the Bayesian model specification and the derivation of important expressions, an in-depth discussion on the rank inference properties of the algorithm, as well as an appendix elaborating how the optimization objective function is derived. We also present additional synthetic data simulations under a variety of scenarios, as well as a set of real-data recommendation and classification experiments to further validate our approach.
I-A Related Work
This paper is primarily concerned with developing a method for joint parameter and model order estimation for PMF tensors. As such, our work is closely related to existing CPD-based PMF tensor estimation approaches [7, 9, 8, 10, 11, 12, 13]. Unlike these works, our method directly estimates the model order (rank) of the PMF tensor from the data, avoiding manual rank selection. Moreover, we take a Bayesian approach tailored to low-rank PMF tensor estimation, and no computation of lower-order marginals is required as in [7, 9, 8, 10, 11].
Nonparametric Bayesian approaches have been used for the probabilistic modeling of categorical random variables [18, 19, 20, 21]. These works use infinite mixtures based on stick-breaking Dirichlet process priors and apply Markov chain Monte Carlo (MCMC) sampling methods for posterior inference. In contrast, we use a finite mixture model, based on simple Dirichlet priors, whose fixed latent structure is a nonnegative CPD which admits a naïve Bayes model interpretation, allowing a universal representation of any joint PMF as well as uniqueness under mild conditions [7]. Although the models in [18] and [20] can be viewed as a CPD, the connection to the naïve Bayes model becomes loose due to the use of infinite mixtures. On the other hand, the models in [19] and [21] admit a higher-order singular value decomposition (HOSVD), which is not unique [4]. Moreover, we adopt variational inference which yields closed-form posterior updates without sampling, unlike MCMC techniques whose convergence can be slow [22].
The CPD representation of a joint PMF tensor, along with the naïve Bayes interpretation, results in a latent variable model that, at first glance, resembles the mixture-of-unigrams model used in document analysis [23, 24]. The mixture-of-unigrams model describes the distribution of words in a single document by selecting one topic (the latent class) and sampling all word tokens independently from that topic, making it a latent class model for repeated observations of a single variable (the document). This differs fundamentally from our model, which directly parametrizes and estimates the full joint PMF of multiple discrete variables. Moreover, our approach also estimates the model order along with the model parameters, whereas the aforementioned works focus on estimating the model parameters.
The automatic rank estimation property of our method relies on a sparsity-promoting Dirichlet prior distribution. Thus, our work can be considered part of the general framework of sparse Bayesian learning, in which Bayesian shrinkage priors play an important role (see, e.g., [25, 26] for an overview). Such shrinkage priors have been used for rank estimation in Bayesian tensor decomposition [27, 28, 29, 30], where the goal is to decompose multidimensional data to discover underlying structures and patterns. Dirichlet process priors have been used in nonparametric Bayesian tensor decomposition, e.g., [31, 32]. In contrast, we use finite Dirichlet priors and our goal is to model and estimate the joint PMF tensor of a set of discrete random variables given a (two-dimensional) dataset of observations. Furthermore, our choice of the Dirichlet prior is informed by the fact that the commonly used Bayesian shrinkage priors (such as the Gaussian scale mixture family [25]) are not suitable for PMF tensor estimation since they do not naturally fulfill the probability simplex constraints.
I-B Notation
@@
This paper uses bold lowercase letters (e.g., $\bm{a}$) to denote vectors, bold uppercase letters (e.g., $\bm{A}$) to represent matrices, and bold calligraphic letters (e.g., $\bm{\mathcal{A}}$) to represent tensors. Uppercase letters (e.g., $A$) denote scalar random variables while lowercase letters (e.g. $a$) denote realizations of the corresponding random variables. The outer product is represented by $\circ$ while T denotes the transpose operator. The squared Frobenius norm is denoted by the operator $\|\cdot\|_{\mathsf{F}}^{2}$.
@@ II Problem Statement
Consider a discrete random vector $\bm{x}=[X_{1},\dotsc,X_{N}]^{\mathsf{T}}\in\mathbb{R}^{N}$ where each random variable $X_{n}$ can take discrete values in $\{1,\dotsc,I_{n}\}$, $n=1,\dotsc,N$. The joint PMF $p(\bm{x})$ can be described by an $N$-dimensional tensor $\bm{\mathcal{X}}\in\mathbb{R}^{I_{1}\times\dots\times I_{N}}$ whereby each element of the tensor is the joint probability of a particular realization of the random variables, i.e., $\mathsf{Pr}(X_{1}=i_{1},\dotsc,X_{N}=i_{N})=\bm{\mathcal{X}}(i_{1},\dotsc,i_{N})$. In general, $\bm{\mathcal{X}}$ can be expressed as a sum of rank-one nonnegative terms via the (nonnegative) CPD
@@
where $\bm{A}_{n}\in\mathbb{R}^{I_{n}\times R}$ denotes a factor matrix with nonnegative elements, and $\bm{A}_{n}(:,r)$ denotes the $r$-th column of $\bm{A}_{n}$. Without loss of generality, the columns of the factor matrices can be restricted to have unit norm such that $\|\bm{A}_{n}(:,r)\|_{p}=1$ for $p\geq 1$, $\forall~n,r$. Then, $\bm{\lambda}=[\lambda_{1},\dotsc,\lambda_{r}]^{\mathsf{T}}$ is the loading vector, consisting of positive elements, which ‘absorbs’ the norms of the columns. The minimum number $R$ of rank-1 terms required for the decomposition to hold is referred to as the rank111or the positive rank [citation] of the tensor. In terms of the CPD, a particular element of the tensor is given by
@@
It has been shown in [6] and [7] that a (nonnegative) rank-$R$ CPD can be interpreted as a particular kind of latent variable model known as the naïve Bayes model. This model assumes that the random variables $\{X_{n}\}_{n=1}^{N}$ are conditionally independent given a hidden (latent) variable $H$ which can take a finite number of $R$ states. Under this interpretation, each element $\lambda_{r}$ of the loading vector is the probability $\mathsf{Pr}(H=r)$ of the hidden variable, while each factor matrix column $\bm{A}_{n}(:,r)$ represents the conditional PMF $p(X_{n}\,|\,H=r)$. Thus, (1) is subject to a set of probability simplex constraints, i.e., $\bm{\lambda}>\bm{0},\bm{A}_{n}\geq\bm{0}$ (nonnegativity) and $\bm{1}^{\mathsf{T}}\bm{\lambda}=1$, $\bm{1}^{\mathsf{T}}\bm{A}_{n}=\bm{1}^{\mathsf{T}}$ (sum-to-one).
@@
(independently for each $n$ and independently of the values in $\bm{x}$), where the outage probability $p$ denotes the probability that $X_{n}$ is unobserved in $\bm{y}$. We have at hand a dataset $\bm{Y}=[\bm{y}_{1},\dotsc,\bm{y}_{T}]$ containing $T$ i.i.d. realizations of $\bm{y}$, where $\bm{y}_{t}=[y_{1,t},\dotsc,y_{N,t}]^{\mathsf{T}}$. The goal is to estimate the CPD components $\{\{\bm{A}_{n}\}_{n=1}^{N},\bm{\lambda}\}$ of the unknown joint PMF tensor $\bm{\mathcal{X}}$ from the observations $\bm{Y}$.